OpenStudy (anonymous):

5. What is the simplest form of the number?

OpenStudy (anonymous):

\[27^{\frac{ 2 }{ 3 }}\]

OpenStudy (anonymous):

A.3 B.6 C.9 D.10

OpenStudy (solomonzelman):

\[\huge\color{blue}{a^{\frac{b}{c}}=\sqrt[c]{a^b}}\]

OpenStudy (solomonzelman):

\[\huge\color{blue}{27^{\frac{2}{3}}=\sqrt[3]{27^2}=9}\]

OpenStudy (solomonzelman):

CLEAR?

OpenStudy (solomonzelman):

Sorry for not coming, I was always either disconnected or busy.

OpenStudy (anonymous):

yes that was good ill be back in a second emergency sorrry

OpenStudy (solomonzelman):

JUST TO SHOW YOU IT IN SIMPLER WAY \[\huge\color{blue}{\sqrt[3]{27^2}=\sqrt[3]{(3^3)^2}=\sqrt[3]{(3^2)^3}=3^2=9}\] do you see the manipulations?

OpenStudy (solomonzelman):

the emergency is fine.